OpenStudy (vshiroky):

Write an equation of the line passing through the points. (4,-16) and (-4,8) The equation of the line is.....

OpenStudy (anonymous):

\[ slope=\frac{8+16}{-4-4}=-3 \]

OpenStudy (anonymous):

(4,-16) and (-4,8) y-8=-3(x+4)=-3x -12 y=-3x -4